# Reference.bib

### v1

@@ 1. Introduction
1.1. Motivation from Griffiths
Hodge-theoretic methods are fundamental in moduli theory of algebraic varieties. Let $\mathcal{X}\rightarrow B$ be a family of smooth projective varieties over some quasi-projective base manifold $B$. It gives a canonical integral variation of Hodge structure over $B$ which induces a period map $\Phi$ from $B$ to some analytic Hodge variety $\Gamma\backslash D$. One may obtain ample information about the family $\mathcal{X}\rightarrow B$ by studying its associated period map $\Phi$. Therefore, the study of local and global behavior of $\Phi$ including its asymptotic behavior near the boundary of $B$ becomes one of the central topics in Hodge theory and algebraic geometry.
@@ 1.1. Motivation from Griffiths
[…]
Griffiths’ problem is almost perfectly addressed when the Hodge variety $\Gamma\backslash D$ is locally Hermitian symmetric. Related works include [BB66] and [AMRT10]. Though these works on the compactifications of locally symmetric varieties are developed without Hodge-theoretic terms, they can be directly applied to many problems in Hodge theory.
However, for the most general cases, a natural completion of period mappings adhere to Griffiths’ problem is still in mystery. There are many attempts to this question emerging in recent years, for example, [GGLR20], [GGR21], [KU08], [KNU10].
@@ 1.2. The main result
The purpose of this paper is to provide a (partial) solution to Griffiths’ long-standing problem. The following informally-stated theorem is the main result of this paper and will be formalized as Theorem 2.8.
@@
[…]
The Conjecture 2.7 will be fully proved in the author’s incoming work [Den25]. In this paper, we will focus on the cases in which the conjecture is evidently true and will result in applications useful in practice.
@@ 1.3. Structure of the paper
In Section 2 we will go over all of the main ingredients in this paper. Necessary backgrounds in Hodge theory will be reviewed in Section 3. In Section 4 we will introduce the concept of combinatorial monodromy complex (CMCX) which is crucial for our construction. Section 5 will be devoted to the construction of the extended period map as a topological morphism. This morphism will be upgraded to an analytic morphism in Section 6. In Section 7 we will discuss applications including some examples in which our construction results in completions in the projective category. In Section 8 we will provide an alternative proof on the algebraicity result for weakly special subvarieties.
Acknowledgment. The author is in debt to Colleen Robles and Matt Kerr for their generous share of ideas. The author also thanks a lot of people for related helpful discussions, including but not limit to: Harold Blum, Robert Friedman, Patricio Gallardo, Phillip Griffiths, Leo Herr, Chikara Nakayama, Christian Schnell, Salim Tayou. A significant part of this work was done during the author’s postdoctoral appointment at Duke University. The author wants to thank the university for the support and environment provided.
@@ 2.1. Overview
$\Phi:B\rightarrow\Gamma\backslash D$
where $\Gamma$ is the monodromy group and $D$ is the corresponding period domain parametrizing polarized Hodge structures of weight $l$ and Hodge numbers $\{h^{p,q}\}_{p+q=l}$. Let $\mathbb{G}:=\mathrm{Aut(D)^{+}}$ which is an algebraic group defined over $\mathbb{Q}$. Let $\mathfrak{g}=\mathrm{Lie}(\mathbb{G})$.
@@ 2.1.1. The classical cases
The locally symmetric variety $\Gamma\backslash D$ has many well-known compactifications, for example, the Satake-Baily-Borel (SBB) compactification [BB66] and the toroidal compactification [AMRT10] which can be described via Hodge-theoretic terms.
Theorem 2.1 ([BB66, AMRT10]).
Given a locally symmetric variety $\Gamma\backslash D$, there are canonical Satake-Baily-Borel compactification $\overline{\Gamma\backslash D}^{\mathrm{SBB}}$ and toroidal compactifications $\overline{\Gamma\backslash D}^{\Sigma}$ depending on the choice of a toroidal fan $\Sigma$, both in the category of projective varieties. Moreover, there is a reduction morphism $\overline{\Gamma\backslash D}^{\Sigma}\rightarrow\overline{\Gamma\backslash D}^{\mathrm{SBB}}$ which may be regarded as partial desingularization.
The SBB and toroidal compactifications are important in the study of moduli spaces of certain geometric objects, especially their degenerations. One of the typical questions in the study of moduli theory is to ask whether a given period mapping $\mathcal{M}\rightarrow\Gamma\backslash D$ over some family of smooth projective varieties over $\mathcal{M}$ can be extended to some compactification of $\mathcal{M}$ with the target space to be the SBB or toroidal compactification of $\Gamma\backslash D$.
In the curve case where $l=1$, let $\Gamma=\mathrm{Sp}_{2g}(\mathbb{Z})$, the period map $\Phi_{g}:\mathcal{M}_{g}\rightarrow\mathfrak{A}_{g}:=\Gamma\backslash D$ associated to the universal genus $g$ curve $\mathcal{M}_{g,1}\rightarrow\mathcal{M}_{g}$ is the Torelli map sending a smooth genus $g$ algebraic curve to its Jacobian. The image of $\Phi_{g}$ is known as the Jacobian locus inside $\mathfrak{A}_{g}$, the moduli space of rank $g$ principal polarized abelian varieties.
@@ 2.1.2. Period mappings of general types
[GRT14] If $D$ is non-classical [^1] The definition of “classical cases” in [GRT14] is a bit different from the definition we use here. and $\Gamma\leq\mathrm{Aut}(D)^{+}$ is an arithmetic subgroup, then the Hodge variety $\Gamma\backslash D$ has no algebraic structure compatible with the analytic structure on $D$.
That is to say, it is hopeless to expect the period map $\Phi:B\rightarrow\Gamma\backslash D$ to be completed in the category of algebraic varieties as in the classical cases. To tackle this difficulty, Kato–Nakayama–Usui (KNU) introduce the theory of space of nilpotent orbits [KU08], [KNU10]. Their theory is a combination of classical asymptotic Hodge theory (e.g. [Sch73], [CKS86]) and the theory of logarithmic geometry. In informal language, one of their main results is:
@@
Assuming a $\Gamma$-strongly compatible weak fan $\Sigma$ exists, the period map $\Phi:B\rightarrow\Gamma\backslash D$ may be completed to a morphism $\Phi:\overline{B}\rightarrow\Gamma\backslash D_{\Sigma}$ in the category of logarithmic manifolds.
@@
  • When $D$ is Hermitian symmetric, we may choose $\Sigma$ in a way such that $\Gamma\backslash D_{\Sigma}$ is exactly the toroidal compactification constructed by [AMRT10]. See [Hay14] for a comparison between these two compactifications in the Siegel space case.
@@
[DR23] When $B$ is a quasi-projective surface, there exists a projective completion $\overline{B}(\Sigma)$ and a KNU space $\Gamma\backslash D_{\Sigma}$ such that the period map $\Phi:B\rightarrow\Gamma\backslash D$ can be completed to $\overline{\Phi}_{\Sigma}:\overline{B}(\Sigma)\rightarrow\Gamma\backslash D_{\Sigma}$ in the category of logarithmic manifolds, and $\mathrm{Img}(\overline{\Phi}_{\Sigma})$ admits a structure of compact complex algebraic space.
@@ 2.2. Main results
More precisely, given a period map $\Phi:B\rightarrow\Gamma\backslash D$, one may factor the period map $\Phi$ via the diagram:
(2.1) $B\xrightarrow{\Phi^{\prime}}\wp^{\prime}\rightarrow\wp:=\mathrm{Img}(\Phi)\hookrightarrow\Gamma\backslash D$
where $B\xrightarrow{\Phi^{\prime}}\wp^{\prime}$ is the Stein factorization of $\Phi$ and is the analytification of a morphism in the category of quasi-projective varieties [BBT22]. Instead of finding a completion of $\Phi:B\rightarrow\Gamma\backslash D$ which requires an enlarged space of the entire Hodge variety $\Gamma\backslash D$, we will complete $\Phi^{\prime}:B\rightarrow\wp^{\prime}$ instead.
@@
CMCX is an analog of a toroidal fan used by [AMRT10] or a $\Gamma$-strongly compatible weak fan used by [KU08] and [KNU10] which does not require any compatibility conditions with the global monodromy group $\Gamma$.
The construction of a CMCX depends on both the $\mathbb{Z}$-polarized variation of Hodge structures and a choice of projective completion $\hat{B}$ of $B$ with simple normal crossing (SNC) divisors. In this sense, we conjecture that:
Conjecture 2.7.
@@
In an upcoming work [Den25], we will show this conjecture is actually a theorem. In this paper we accept this conjecture as an assumption. In Section 4 we will see that under some mild assumptions, the conjecture is true and the construction of the CMCX is evident.
@@
Let $B\xrightarrow{\Phi^{\prime}}\wp^{\prime}\rightarrow\wp$ be the Stein factorization of $B\xrightarrow{\Phi}\wp$ and $\hat{B}$ be a fixed completion of $B$ with SNC boundary divisors. Let $\mathcal{T}$ be a combinatorial monodromy complex associated to $\hat{B}$. There exists a completion of $\overline{B}$ of $B$, a completion $\overline{\wp}$ of $\wp^{\prime}$, such that we have the following extension in the category of complex analytic spaces:
[…]
@@
Indeed, the nilpotent orbit $(\tau_{b^{\prime}},F_{b^{\prime}})$ in Theorem 2.8 is determined by the classical theory of the nilpotent orbit theorem by Schmid. See Section 3 for an overview.
The proof of Theorem 2.8 will be done in two steps: In Section 5 we construct the map $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ on the set-theoretic level, this is Theorem 5.1. In Section 6 we will show $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ has a natural structure in the category of complex analytic spaces.
In the most general cases, it is unrealistic to describe the construction of $\overline{\Phi}$ and the completion $\overline{B}$ of $B$. However, under some certain assumptions the constructions are elementary:
@@
Motivated from the construction of [AMRT10] and the quasi-projectivity of period images [BBT22], it is natural to conjecture:
@@
The construction in Theorem 2.8 can always be done in a way such that $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ lies in the category of projective varieties.
@@ 2.3.1. Comparison with [DR23]
given by Theorem 2.5 in [DR23] is defined on some logarithmic modification $\overline{B}(\Sigma)$ of $\overline{B}$. It is in general not possible to describe the weak fan $\Sigma$ or the projective surface $\overline{B}(\Sigma)$ over which the map $\overline{\Phi}_{\Sigma}$ lives in.
@@
which does not require the existence of any extension of the entire Hodge variety $\Gamma\backslash D$. In this case, the completion $\overline{B}$ can be any smooth completion of $B$ with $\overline{B}-B$ a normal crossing divisor.
As a summary, though the main theorem of this paper does not provide a completed Hodge variety $\Gamma\backslash D_{\Sigma}$ which plays an important role in KNU’s theory [KU08], [KNU10], the completion $\overline{B}$ over which the extended period map exists is much easier to describe.
2.3.2. Comparison with [DGGR25]
In [DGGR25], Mark Green, Phillip Griffiths, Colleen Robles and the author will give a similar construction. It is known by the Schmid’s nilpotent orbit theorem (Theorem 3.3) that for any $b\in\overline{B}$, there exists a unique $\Gamma$-class of nilpotent orbits $(\sigma_{b},F_{b})$ associated to $b\in\overline{B}$ via the asymptotic behavior of $\Phi$ near $b$. Instead of trying to record the information on both the nilpotent cone $\sigma_{b}$ and the limiting Hodge flag $F_{b}$, we only try to have $\overline{\Phi}$ encode the nilpotent orbit $\exp(\sigma_{b,\mathbb{C}})F_{b}$. In other words, the information on the submonoid $\sigma_{b}\subset\sigma_{b,\mathbb{C}}$ is lost.
The advantage of the construction in [DGGR25] is that we do not need a combinatorial monodromy complex which encodes the information about monodromy cones. This makes the construction simpler. However, in this case we are only able to show $\overline{\wp}$ is a Hausdorff topological space which admits a stratification by complex analytic spaces in general, and $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ is only a topological completion.

### v4

# On the generalized toroidal completion of period mappings

@@ 1. Introduction
1.1. Motivation
Hodge-theoretic methods are fundamental in moduli theory of algebraic varieties. Let $\mathcal{X}\rightarrow B$ be a family of smooth projective varieties over some quasi-projective base manifold $B$. It gives a canonical integral variation of Hodge structure over $B$ which induces a period map $\Phi$ from $B$ to some analytic Hodge variety $\D\Gamma\backslash$ D. One may obtain ample information about the family $\mathcal{X}\rightarrow B$ by studying its associated period map $\Phi$. Therefore, the study of local and global behavior of $\Phi$ including its asymptotic behavior near the boundary of $B$ becomes one of the central topics in Hodge theory and algebraic geometry.
@@ 1.1. Motivation
    When $\Phi$ is the period map underlying a family of smooth projective varieties $\mathcal{X}\xrightarrow{\pi}B$, we may study singular objects along $\overline{B}-B$ by looking at the boundary limiting Hodge data from $\overline{\Phi}$.
Griffiths’ problem is almost perfectly addressed when the Hodge variety $\D\Gamma\backslash$ D is locally Hermitian symmetric. Related works include [BB66] and [Ash+10]. Though these works on the compactifications of locally symmetric varieties are developed without Hodge-theoretic terms, they can be directly applied to many problems in Hodge theory.
For the most general cases, an analogue of the Baily-Borel compactification was recently announced in this generality in [Bak+25], following earlier works in [BBT22],[Gre+23],[GGR21]. In this paper, we answer this question in a different direction by constructing analogues of toroidal compactifications following [KU08], [KNU10].
@@ 1.2. The main result
The purpose of this paper is to provide a solution to Griffiths’ long-standing problem. The following informally-stated theorem is the main result of this paper and will be formalized as Theorem 2.8.
@@
Denote $\D\mathrm{Im}(\Phi)=:\wp\subset\Gamma\backslash$ D and let $\Phi^{\prime}:B\rightarrow\wp^{\prime}$ be the Stein factorization of $\Phi:B\rightarrow\wp$. There exists a projective completion $\overline{B}$ of $B$, a completion $\overline{\wp}$ of $\wp^{\prime}$ as complex analytic spaces, such that $\Phi$ may be completed as a morphism of compact complex analytic spaces $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$. Moreover, up to finite data, $\overline{\Phi}(b)$ records the limiting mixed Hodge structures at $b\in\overline{B}$.
@@
As a consequence, the main theorem of this paper may be regarded as a generalization of [Ash+10, Chap. 1-3] to any period mappings defined over some quasi-projective variety. It is natural to ask whether the projectivity result in [Ash+10, Chap. 4] also generalizes (Conjecture 2.12).
@@ 1.3. Structure of the paper
In Section 2 we will go over all of the main ingredients in this paper. Necessary backgrounds in Hodge theory will be reviewed in Section 3. In Section 4 we will introduce the concept of combinatorial monodromy complex (CMCX) which is crucial for our construction. Section 5 will be devoted to the construction of the extended period map as an analytic morphism. In Section 6 we will discuss applications including some examples in which our construction results in completions in the projective category. In Section 7 we will provide an alternative proof on the algebraicity result for weakly special subvarieties. Finally in Section 8 and 9, we will prove the general existence of CMCX.
Acknowledgment. The authors thank a lot of people for related helpful discussions, including but not limit to: Ben Bakker, Harold Blum, Robert Friedman, Patricio Gallardo, Phillip Griffiths, Leo Herr, Matt Kerr, Chikara Nakayama, Colleen Robles, Christian Schnell, Salim Tayou. The first author wants to thank Duke University for great environment provided to postdoctoral scholars. The second author wants to thank the Institute for Advanced Study and the Ashvin B. Chhabra and Daniela Bonafede-Chhabra Membership in particular for their support.
@@ 2.1. Overview
$\D\Phi:B\rightarrow\Gamma\backslash$ D
where $\Gamma$ is the monodromy group and $D$ is the corresponding period domain parametrizing polarized Hodge structures of weight $l$ and Hodge numbers $\{h^{p,q}\}_{p+q=l}$. Let $\mathbf{G}:=\mathrm{Aut(D)^{+}}$ which is an algebraic group defined over $\mathbb{Q}$. Let $\mathfrak{g}=\mathrm{Lie}(\mathbf{G})$.
@@ 2.1.1. The classical cases
The locally symmetric variety $\D\Gamma\backslash$ D has many well-known compactifications, for example, the Satake-Baily-Borel (SBB) compactification [BB66] and the toroidal compactification [Ash+10] which can be described via Hodge-theoretic terms.
Theorem 2.1 ([BB66, Ash+10]).
Given a locally symmetric variety $\D\Gamma\backslash$ D, there are canonical Satake-Baily-Borel compactification $\overline{\Gamma\backslash D}^{\mathrm{SBB}}$ and toroidal compactifications $\overline{\Gamma\backslash D}^{\Sigma}$ depending on the choice of a toroidal fan $\Sigma$, both in the category of projective varieties. Moreover, there is a reduction morphism $\overline{\Gamma\backslash D}^{\Sigma}\rightarrow\overline{\Gamma\backslash D}^{\mathrm{SBB}}$ which may be regarded as partial desingularization.
The SBB and toroidal compactifications are important in the study of moduli spaces of certain geometric objects, especially their degenerations. One of the typical questions in the study of moduli theory is to ask whether a given period mapping $\D\mathcal{M}\rightarrow\Gamma\backslash$ D over some family of smooth projective varieties over $\mathcal{M}$ can be extended to some compactification of $\mathcal{M}$ with the target space to be the SBB or toroidal compactification of $\D\Gamma\backslash$ D.
In the curve case where $l=1$, let $\Gamma=\mathrm{Sp}_{2g}(\mathbb{Z})$, the period map $\D\Phi_{g}:\mathcal{M}_{g}\rightarrow\mathfrak{A}_{g}:=\Gamma\backslash$ D associated to the universal genus $g$ curve $\mathcal{M}_{g,1}\rightarrow\mathcal{M}_{g}$ is the Torelli map sending a smooth genus $g$ algebraic curve to its Jacobian. The image of $\Phi_{g}$ is known as the Jacobian locus inside $\mathfrak{A}_{g}$, the moduli space of rank $g$ principal polarized abelian varieties.
@@ 2.1.2. Period mappings of general types
[GRT14] If $D$ is non-classical [^1] The definition of “classical cases” in [GRT14] is a bit different from the definition we use here. and $\Gamma\leq\mathrm{Aut}(D)^{+}$ is an arithmetic subgroup, then the Hodge variety $\D\Gamma\backslash$ D has no algebraic structure compatible with the analytic structure on $D$.
That is to say, it is hopeless to expect the period map $\D\Phi:B\rightarrow\Gamma\backslash$ D to be completed in the category of algebraic varieties as in the classical cases. To tackle this difficulty, Kato–Nakayama–Usui (KNU) introduce the theory of space of nilpotent orbits [KU08], [KNU10]. Their theory is a combination of classical asymptotic Hodge theory (e.g. [Sch73], [CKS86]) and the theory of logarithmic geometry. In informal language, one of their main results is:
@@
Assuming a $\Gamma$-strongly compatible weak fan $\Sigma$ exists, the period map $\D\Phi:B\rightarrow\Gamma\backslash$ D may be completed to a morphism $\Phi:\overline{B}\rightarrow\Gamma\backslash D_{\Sigma}$ in the category of logarithmic manifolds.
@@
  • When $D$ is Hermitian symmetric, we may choose $\Sigma$ in a way such that $\Gamma\backslash D_{\Sigma}$ is exactly the toroidal compactification constructed by [Ash+10]. See [Hay14] for a comparison between these two compactifications in the Siegel space case.
@@
[DR23] When $B$ is a quasi-projective surface, there exists a projective completion $\overline{B}(\Sigma)$ and a KNU space $\Gamma\backslash D_{\Sigma}$ such that the period map $\D\Phi:B\rightarrow\Gamma\backslash$ D can be completed to $\overline{\Phi}_{\Sigma}:\overline{B}(\Sigma)\rightarrow\Gamma\backslash D_{\Sigma}$ in the category of logarithmic manifolds, and $\mathrm{Img}(\overline{\Phi}_{\Sigma})$ admits a structure of compact complex algebraic space.
@@ 2.2. Main results
More precisely, given a period map $\D\Phi:B\rightarrow\Gamma\backslash$ D, one may factor the period map $\Phi$ via the diagram:
[…]
where $B\xrightarrow{\Phi^{\prime}}\wp^{\prime}$ is the Stein factorization of $\Phi$ and is the analytification of a morphism in the category of quasi-projective varieties [BBT22]. Instead of finding a completion of $\D\Phi:B\rightarrow\Gamma\backslash$ D which requires an enlarged space of the entire Hodge variety $\D\Gamma\backslash$ D, we will complete $\Phi^{\prime}:B\rightarrow\wp^{\prime}$ instead.
@@
CMCX is an analog of a toroidal fan used by [Ash+10] or a $\Gamma$-strongly compatible weak fan used by [KU08] and [KNU10] which does not require any compatibility conditions with the global monodromy group $\Gamma$.
The construction of a CMCX depends on both the $\mathbb{Z}$-polarized variation of Hodge structures and a choice of projective completion $\hat{B}$ of $B$ with simple normal crossing (SNC) divisors. Assuming both ingredients, we have:
Theorem 2.7.
@@
Theorem 2.7 will be proved in general in Section 9. The techniques are purely combinatorial. In Section 4 we will see that under some mild assumptions, the construction of the CMCX is evident.
@@
Let $B\xrightarrow{\Phi^{\prime}}\wp^{\prime}\rightarrow\wp$ be the Stein factorization of $B\xrightarrow{\Phi}\wp$ and $\hat{B}$ be a fixed completion of $B$ with SNC boundary divisors. Let $\mathcal{T}$ be a combinatorial monodromy complex associated to $\hat{B}$. There exists a completion $\overline{B}$ of $B$ compatible with $\mathcal{T}$, a completion $\overline{\wp}$ of $\wp^{\prime}$, such that we have the following extension in the category of complex analytic spaces (and consequently, algebraic spaces):
(2.2) ${\wp^{\prime}}B¯{\overline{B}}$$\overline{\wp}$$\scriptstyle{\Phi^{\prime}}$$\scriptstyle{\overline{\Phi}}$
@@
Indeed, the nilpotent orbit $(\tau_{b^{\prime}},F_{b^{\prime}})$ in Theorem 2.8 is determined by the classical theory of the nilpotent orbit theorem by Schmid in [Sch73]. See Section 3 for an overview.
Remark 2.10.
In particular, this gives a new proof of the algebraicity of period images, proven in [BBT22], up to normalization.
The proof of Theorem 2.8 will be done in two steps: In Section 5 we construct the map $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ on the set-theoretic level and show $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ has a natural structure in the category of complex analytic spaces. In the most general cases, describing the construction of $\overline{\Phi}$ and the completion $\overline{B}$ of $B$ could be complicated on the computational side. However, under some certain assumptions the constructions are elementary:
@@
Motivated from the construction of [Ash+10] and the quasi-projectivity of period images [BBT22], it is natural to conjecture:
@@
The construction in Theorem 2.8 can always be done in a way such that $\overline{\wp}$ is a projective variety.
@@ 2.3. Applications and related works
We introduce some applications of our main result as well as its interaction with several related works.
@@ 2.3.1. Comparison with [DR23]
given by Theorem 2.5 in [DR23] is defined on some logarithmic modification $\overline{B}(\Sigma)$ of $\overline{B}$. Due to the fact that the global monodromy group $\Gamma$ is mostly immeasurable, it is in general not possible to describe the weak fan $\Sigma$ or the projective surface $\overline{B}(\Sigma)$ over which the map $\overline{\Phi}_{\Sigma}$ lives in.
@@
which does not require the existence of any extension of the entire Hodge variety $\D\Gamma\backslash$ D. In this case, the completion $\overline{B}$ can be any smooth completion of $B$ with $\overline{B}-B$ a normal crossing divisor.
As a summary, though the main theorem of this paper does not provide a completed Hodge variety $\Gamma\backslash D_{\Sigma}$ which plays an important role in KNU’s theory [KU08], [KNU10], in many cases the completion $\overline{B}$ over which the extended (Stein-factorized) period map exists is much easier to describe.
2.3.2. Relation with [Bak+25]
In this paper the authors construct a canonical completion $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ of $\Phi:B\rightarrow\wp$. The construction is called as the generalized (Satake)–Baily–Borel compactification, as the “hodge-theoretic data” that is recorded at the boundary is only that of pure hodge structures (the graded components) with no extension data. It recovers the classical Satake–Baily–Borel compactification in the case when the period domain $D$ is Hermitian symmetric. As an analog of the relation between Satake–Baily–Borel compactification and toroidal compactification in the classical case, studying the relationship between the generalized completions is a potentially interesting topic.
In [Bak+25] the authors compactify the actual period image, without needing to pass to the Stein factorization. It seems very plausible the same thing is possible with out main theorem.